OpenStudy (anonymous):

if the geometric mean of two numbers is 6 and its arithmetic mean is 10 find the numbers.!!

OpenStudy (anonymous):

let the numbers be x and y, then x+y=2*10 x*y=6^2 solve for x and y

OpenStudy (anonymous):

Solve it for me please.

OpenStudy (anonymous):

here will come a quadratic equation if we put x=36/y in equation 1

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

\[\frac{ 36 }{ y }+y=20\] \[y^2-20y+36=0\] \[y^2-2y-18y+36=0 \] now you can solve.

OpenStudy (anonymous):

then it will give us 2 value for y.. i want one.

OpenStudy (anonymous):

when y=2 ,x=36/2=18 when y=18, x=36/2 =18 so only order is different. hence numbers are 2 and 18